OpenStudy (xapproachesinfinity):

maximization problem Build a rectangle pen with three parallel partitions using 400 feet fencing. what dimensions will maximize the the total area.

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

Checking if I'm right! we have P=2y+4x=400 y=200-2x A=xy ====> A=x(200-2x) A'=200-4x ===>200-4x=0 ===> x=50 y=200-100=100

OpenStudy (xapproachesinfinity):

I'm not sure about three parallel partitions thingy! is the diagram correct

OpenStudy (amistre64):

the diagram is fine

OpenStudy (amistre64):

f = xy g = 4x + 2y - 400 using lagrange multiplier method: when the gradient of f and g are simply a multiple of each other, then we have critical points to play with fx = y Lgx = 4L fy = x Lgy = 2L x = 2L y = 4L by substitution: g = 0 = 4(2L) + 2(4L) - 400 400 = 16L, L = 25 therefore x = 50, and y = 100 :)

OpenStudy (xapproachesinfinity):

oh seems i was right! i have no idea about Lagrange multiplier method our professor didn't mention that in the course

OpenStudy (amistre64):

didnt mention it in my course either ... had to go figure it out all on me own ;)

OpenStudy (xapproachesinfinity):

oh i see! good thing any page, book to check that method i looks nice lol

OpenStudy (amistre64):

none that come to mind off hand. pauls calculus site might be a good primer in it.

OpenStudy (xapproachesinfinity):

ok thanks, i will check that! =================== just a question came to my mind, when i solved couple of optimization problem i find solutions when i check my answer, i find that they use some boundaries like this example: an open rectangle box with square base is to be constructed from 48 ft^2 of material. What dimensions will give the largest possible volume when they solved they said 0<x<= sqrt(48) i know that x cannot be zero how about sqrt part?

OpenStudy (xapproachesinfinity):

I found solution without those boundaries though

OpenStudy (amistre64):

if we have a function and multiple constaints such as say: f, constrained by g and h then the gradient of f is equal to a linear combination of the gradient of the constraints. \[\nabla f=\lambda \nabla g+ \kappa \nabla h\]

OpenStudy (amistre64):

an open rectangle box with square base is to be constructed from 48 ft^2 of material. What dimensions will give the largest possible volume when they solved they said 0<x<= sqrt(48) ------------------------------------ now the question is: do we use all 48 ft^2 or do we assume its a square shape of material and that there is cut offs?